### v2

# Learning to Extract Rational Evidence via Reinforcement Learning for Retrieval-Augmented Generation

Abstract
Retrieval-Augmented Generation (RAG) effectively improves the accuracy of Large Language Models (LLMs). However, retrieval noises significantly impact the quality of LLMs’ generation, necessitating the development of denoising mechanisms. Previous methods extract evidence straightforwardly without explicit thinking, which risks filtering out key clues and struggles with generalization. To this end, we propose LEAR, which learns to extract rational evidence by (1) explicitly reasoning to identify potential cues within retrieval contents first, and then (2) consciously extracting to avoid omitting any key cues helpful for answering questions. Specifically, we frame evidence reasoning and extraction into one unified response for end-to-end training; apply knowledge token masks for disentanglement to derive reasoning-based and extraction-based answers; and devise three types of verifiable reward functions, including answer, length, and format, to update the model via the policy optimization algorithm. Extensive experiments on three benchmark datasets show the effectiveness of LEAR, providing compact and high-quality evidence, improving the accuracy of downstream tasks, and promoting effective application in online RAG systems [^1] The code, data, and models will be available at https://github.com/HITsz-TMG/LEAR..
Figure 1: Motivating example, where key clues are marked in green: 1 The key clues are filtered out, and LLMs answer incorrectly; 2 The key clues are extracted successfully, guided by the evidence reasoning.
1 Introduction
Retrieval-Augmented Generation (RAG) prevails in Large Language Models (LLMs). It has been shown highly effective for many knowledge-intensive tasks Lewis et al. (2020); Wu et al. (2022), such as open-domain question answering Shi et al. (2024); Trivedi et al. (2023), fact-checking Du et al. (2023); Zhao et al. (2024b), and dialog generation Izacard et al. (2023); Thoppilan et al. (2022), to produce more faithful and reliable outputs. In ideal conditions, LLMs should be grounded on purely relevant retrieval contents to generate accurate output and also facilitate inference. However, due to imperfect retrieval systems or noisy data in the retrieval corpus Wang et al. (2023); Wei et al. (2024); Zhao et al. (2024a), the retrieval contents usually contain lots of irrelevant or noisy snippets, which distract LLMs’ attention and also inflict a heavy blow on the generation quality of LLMs. Therefore, it is necessary and valuable to extract evidence as well as filter out noise for RAG to achieve superior performance.
Recently, several studies have attempted to address this issue. They can be mainly divided into two categories: (1) Reranking more relevant passages go up to the top of the retrieval list Hwang et al. (2024b); Mortaheb et al. (2025); Mao et al. (2021); and (2) Summarizing retrieval contents into a coherent narrative with higher signal to noise ratio Wang et al. (2023); Zhao et al. (2024a); Zhu et al. (2024); Xu et al. (2024). The former heavily relies on the performance of chosen rerankers and may disrupt context after reranking. It usually performs worse than the latter due to a lack of contextual understanding. The latter aims to train LLMs as a filtering model through fine-tuning Wang et al. (2023) or preference optimization Zhao et al. (2024a). They typically create the training data using hand-crafted strategies, e.g., String Inclusion that measures whether the golden answer is included in the candidate passage, and Lexical Overlap, calculating the unigram overlap between the golden answer and candidate passage.
Despite effectiveness, existing methods re-rank or summarize retrieval contents straightforwardly without explicit thinking, which risks filtering out key clues due to a lack of a deep understanding of retrieved contents. Figure 1 shows an example of the evidence extraction (summarization) for question answering: 1 The extracted evidence fails to maintain key clues and results in a wrong answer; and 2 The evidence reasoning guides the following evidence extraction, e.g.,Passage 2 is relevant and mentions……”, where key clues are successfully maintained in the extracted evidence and thus LLMs answer correctly. This motivates us to develop a rational evidence extractor for RAG, which deeply reasons and then consciously extracts as shown in the bottom layer of Figure 1.
In this work, we propose LEAR, which LEArns to extract Rational evidence via reinforcement learning for RAG. Specifically, LEAR frames evidence reasoning (enclosed within <reason> and </reason>) and extraction (enclosed within <extract> and </extract>) into one response for end-to-end training and generates multiple responses to collect trajectories that contain good and bad experiences from the policy. After that, LEAR applies two knowledge token masks on each response to disentangle evidence reasoning and extraction for deriving reasoning-based and extraction-based answers (enclosed within <answer> and </answer> tags) to assess their quality respectively. Finally, we devise three types of rule-based verifiable reward functions, including answer, length, and format, to guide model optimization via Group Relative Policy Optimization (GRPO) Shao et al. (2024). As such, LEAR extracts rational evidence by reasoning about retrieval contents and then consolidating key cues into a concise and useful narrative, which is fed into LLMs to generate more accurate output. Our main contributions can be summarized as follows:
  • We unveil the major issue that hinders evidence extraction for RAG, i.e., lacking a deep contextual understanding of retrieval contents, which risks filtering out key clues.
  • We propose a novel rational evidence extraction learning framework, LEAR, which learns to discern potential cues in retrieval contents and then consciously consolidates any key cues into a concise yet very helpful evidence.
  • We conduct extensive experiments on three benchmark datasets, where the results fully demonstrate the superiority of LEAR in terms of performance, generalization, efficiency, as well as robustness against retrieval noise.
2 Preliminaries
2.1 Problem Statement
In RAG Lewis et al. (2020); Wu et al. (2022), LLMs are given an input query $q$ and a set of top-$k$ retrieved passages $P=\{p_{1},p_{2},...,p_{k}\}$, which aim to generate an output $o$ closing to the golden answer $a$, conditioned on ($q$, $P$). In the traditional RAG paradigm, top-$k$ retrieved passages are directly fed into LLMs. However, retrieved passages $P$ often contain irrelevant or noisy contents Wang et al. (2023); Zhao et al. (2024a), which considerably degrade LLMs’ generation quality. Given that, an additional evidence extraction model is proposed to condense $P$ into a concise, query-relevant context $e$ to improve generation quality and speed. Formally, the paradigm of RAG with evidence extraction can be formulated as follows:
$e=\mathcal{M}_{\mathcal{E}}(\cdot|q,P),\quad o=\mathcal{M}_{\mathcal{G}}(\cdot|q,e),$ (1)
where $\mathcal{M}_{\mathcal{E}}(\cdot)$ denotes the evidence extractor; $e$ is the extracted evidence; $\mathcal{M}_{\mathcal{G}}(\cdot)$ denotes the answer generator, i.e., a LLM; and $o$ is the generated output. Despite effectiveness, we argue that this vanilla paradigm may risk filtering out key clues due to a lack of deep reasoning on retrieved passages. Going beyond the previous paradigm, we present a novel one of RAG with rational evidence extraction, which  can  be  formulated  as  follows:
$\begin{split}e&=\mathcal{M}_{\mathcal{E}}(\cdot|q,P,r)\cdot\mathcal{M}_{\mathcal{E}}(r|q,P),\\ o&=\mathcal{M}_{\mathcal{G}}(\cdot|q,e),\end{split}$ (2)
where we train the model in an on-policy manner, meaning that the evidence extractor $\mathcal{M}_{\mathcal{E}}(\cdot)$ and the answer generator $\mathcal{M}_{\mathcal{G}}(\cdot)$ are the same one during training; $r$ denotes evidence reasoning. It ($r$) explicitly and thoroughly identifies any cues lying in retrieved passages ($P$) to guide the evidence extraction ($e$) to  achieve  more  accurate  output ($o$). The overall objective is to learn a policy $\pi_{\theta}:e\sim\mathcal{M}_{\mathcal{E}}(\cdot|q,P,r)\cdot\mathcal{M}_{\mathcal{E}}(r|q,P)$ to extract a rational evidence $e$ that maximizes the expected utility $\mathcal{U}(a,o)$ over the given query distribution $\mathcal{D}$:
$\max_{\pi_{\theta}}~{}$ $\mathbb{E}_{a\sim\mathcal{M}_{\mathcal{G}}(\cdot|q,e),~{}e\sim\pi_{\theta}(q,P),~{}q\sim\mathcal{D}}\Big{[}\mathcal{U}(a,o)\Big{]},$
$s.t.\ |e|\ll|P|$ (3)
where $\mathcal{U}(\cdot)$ is a utility function (e.g., $\mathrm{F}_{1}$ score) and is used to measure the quality of outputs conditioned on the golden answer $a$; $|e|$ and $|P|$ denote the length of evidence  and  passages, respectively.
(a) Results on NQ.
(b) Results on HotoptQA.
Figure 2: The results w.r.t. AR on the in-distribution NQ as  well  as  out-of-distribution  HotoptQA datasets.
Dataset Vanilla Evidence Rational Evidence Rationale
NQ 70.79% 75.24% 77.30%
HotoptQA 60.55% 67.74% 71.48%
Table 1: AR results on Vanilla Evidence, Rational Evidence and Rationale w.r.t. NQ and HotpotQA datasets.
2.2 Empirical Study
We conduct an empirical study to verify the assumption that evidence extractors can retain as many key clues as possible via the paradigm of reasoning first and then extracting. We construct a synthetic instruction dataset by leveraging DeepSeek-R1222https://api-docs.deepseek.com/guides/reasoning_model DeepSeek-AI et al. (2025), where we sample 1K instances from the training set of Natural Question (NQ) Kwiatkowski et al. (2019). The output of each instance in the dataset contains three parts: (1) <reason> evidence reasoning </reason>; (2) <extract> evidence extraction </extract>; and (3) <answer> final answer </answer>. We filter out instances with incorrect final answers, resulting in about 620 instances. Then, we create two variants of the dataset, where the output of the first one consists of “<reason></reason><extract></extract>”; that of the second one only consists of “<extract></extract>”. Finally, we fine-tune Qwen2.5-1.5B-Instruct Yang et al. (2025) on these two synthetic instruction datasets through the LoRA (Low-Rank Adaptation) Hu et al. (2022), respectively, where the LoRA Rank is set to 16; we set the number of epochs as 3. We tested on in-distribution NQ and out-of-distribution HotoptQA Yang et al. (2018) datasets, excluding instances from the test set [^3] Because the test set is not available for HotoptQA, we serve the development set of HotoptQA as the test set., where the retrieved passages did not contain the golden answer. For a fair comparison, we set the maximum length of the extracted evidence as 100.
Table 1 and Figure 2 show the experimental results w.r.t. Answer Recall (AR) Zhao et al. (2025); Jiang et al. (2024), which measures the recall of the golden answer string in the extracted evidence. We denote the evidence extracted by the model trained on the first and second datasets as “rational evidence” (marked as $\mathrm{\mathbb{M}}$) and “vanilla evidence” (marked as $\mathrm{\mathbb{N}}$), respectively. We use “rationale” to denote the evidence reasoning. Taking $\mathrm{\mathbb{M}}\land\neg\mathrm{\mathbb{N}}$ for example, it means the answer string is recalled by the rational evidence but not recalled by the vanilla one. From the results, we have the following observations: (1) The performance with rational evidence is consistently better than that with vanilla one, whether on in-distribution or out-of-distribution datasets. For example, on NQ, rational evidence achieves the AR of 75.24%, while vanilla one achieves that of 70.79%; (2) The performance with rationale is best, manifesting the necessity of reasoning first. The performance with rational evidence is slightly worse than that with rationale, indicating that a better optimization is needed to bridge this gap; and (3) The percentage of $\mathrm{\mathbb{M}}\land\neg\mathrm{\mathbb{N}}$ is considerably higher than that of $\neg\mathrm{\mathbb{M}}\land\mathrm{\mathbb{N}}$, e.g., 14.49% vs. 7.30% on HotpotQA, fully demonstrating the superiority of the rational evidence.
Figure 3: The overall system framework of the proposed LEAR and previous vanilla paradigm. (a) Compared to vanilla evidence, our rational evidence is more skilled in extracting key cues. (b) Existing works learn to extract vanilla evidence via supervised fine-tuning, where the output, i.e., oracle that contains key clues, is usually created by hand-crafted strategies, e.g., String Inclusion and Lexical Overlap in Wang et al. (2023). (c) The proposed framework LEAR incentivizes rational evidence extraction capability in the extractor $\pi_{\theta}$ via reinforcement learning.
We also provide a detailed failure analysis of vanilla evidence in Appendix B. Briefly, the failures can be attributed to “3I”: (1) Incompleteness, where the evidence provides some relevant information but lacks the key clues; (2) Irrelevance, where the evidence provides information about the wrong entity, event, or topic; (3) Inaccuracy, where the evidence contains incorrect information, such as wrong dates and names. In contrast, rational evidence is able to provide accurate and relevant information as well as locate the specific entity or detail needed for the answer. Given the above qualitative as well as quantitative analyses, it is necessary and valuable to learn to extract rational evidence for improving RAG performance.
3 Methodology
The overall framework of LEAR is illustrated in Figure 3. In this section, we first introduce the acquisition of rational evidence (§3.1). Then, we define three verifiable reward functions (§3.2). Finally, we describe the training of the extractor through online policy optimization in detail (§3.3).
3.1 Obtaining Rational Evidence
As stated in §1, vanilla evidence risks filtering out key clues Wang et al. (2023); Zhao et al. (2024a). An empirical study (§2.2) further validates our assumption, where rational evidence is significantly better than vanilla one. Thus, we aim to optimize the evidence extractor $\mathcal{M}_{\mathcal{E}}(\cdot)$ to learn to extract rational evidence for RAG, formulated as: $e\sim\mathcal{M}_{\mathcal{E}}(\cdot|q,P,r)\cdot\mathcal{M}_{\mathcal{E}}(r|q,P)$. Specifically, we feed the query $q$ and its corresponding retrieved passages $P$ into the evidence extractor $\mathcal{M}_{\mathcal{E}}(\cdot)$ and instruct the evidence extractor to reason first and then extract. The evidence reasoning and extraction are enclosed within the special reason and extract tags, respectively, i.e., <reason></reason> and <extract></extract>. The prompt used for rational evidence generation is provided in Appendix A. Given the rationale and the rational evidence, a question naturally arises: How to assess the quality of them? In Section 2.2, we assess the quality of them via answer recall, a heuristic metric. Analogously, previous works, taking Wang et al. (2023) for example, heuristically employ String Inclusion or Lexical Overlap to measure the answer recall or unigram overlap. However, we argue that these metrics are not aligned with the ultimate goal of RAG, i.e., to generate output as close to the  golden  answer  as  possible.
To this end, we propose assessing the quality of the rationale and rational evidence by assessing the generated outputs conditioned on them. It is well known that causal attention mechanisms Radford et al. (2018); Vaswani et al. (2017) in LLMs aggregate information from previous contexts. Given that, another question arises: How to generate the output conditioned on them respectively without information leakage? To achieve the above goal, we apply two knowledge token masks on each response to disentangle the rationale and rational evidence: (1) The first one masks the rational evidence $e$ to assess the quality of the rationale $r$; (2) In contrast, the second one masks the retrieved passages $P$ and rationale $r$ simultaneously to assess the quality of the rational evidence $e$. Here, we adopt hard masking on input rather than soft masking on attention, because soft masking will cause information leakage where the hidden states have already aggregated information from previous contexts. Thus, we prefill input contexts from scratch after masking and generate three  outputs  conditioned  on  different  contexts:
[…]
(a) $y=\frac{1}{1+e^{-x/\tau}}$.
(b) $y=x^{\gamma}$.
Figure 4: Illustration of length rewards in terms of different skewness or smoothness controlled by $\tau$ and $\gamma$, where we directly plot  the  basic  functions  for  clarity.
3.2 Modeling Verifiable Reward
The reward function plays a key role in Reinforcement Learning (RL) Ouyang et al. (2022); Wang et al. (2024); DeepSeek-AI et al. (2025), which guides the optimization process towards desired properties. To train LEAR via RL, we design three types of rule-based verifiable reward functions w.r.t. three primary desired properties for evidence extraction: (1) The correctness of the generated outputs; (2) The comprehensiveness of rationale and the conciseness of rational evidence; and (3) The format specification of the responses. For reward modeling, we first collect a set of Preao-tuple $<P,r,e,a,o_{r,e,f}>$, where a Preao-tuple e is composed of retrieved Passgaes $P$, rationale $r$, rational evidence $e$, golden answer $a$, generated output $o_{r,e,f}$ conditioned on different contexts. Given that, we design three types of verifiable reward functions to assess the rational evidence extractor w.r.t. the three desired properties:
  • Answer Reward. It focuses on the correctness of the final outputs enclosed within <answer> and </answer> tags. However, different downstream tasks of RAG (e.g., QA, fact verification, and dialog generation) use different metrics to evaluate model predictions, including Exact Match (EM), $\mathrm{F}_{1}$ score, and Accuracy. Given that, a question naturally arises: How to measure the answer reward uniformly w.r.t. different downstream tasks of RAG? Here, we employ the unigram $\mathrm{F}_{1}$ to measure the answer rewards in a unified and balanced manner Song et al. (2025):
    $R_{*}^{ans}=\mathrm{F}_{1}(a,o_{*}),\,\,\,*\in\{r,e,f\},$ (5)
    where $R_{*}^{ans}\in[0,1]$ denotes the answer reward for the generated $o_{*}$. If $o_{*}$ is similar to $a$, then $R_{*}^{ans}$ is close to 1.0; otherwise, it is close to 0.0.
  • Length Reward. It focuses on two aspects. The first one is the comprehensiveness of rationale, enclosed within <reason> and </reason> tags, where the rationale usually needs to be relatively long to identify any cues lying in retrieval contents. The second one is the conciseness of rational evidence, enclosed within <extract> and </extract> tags, where the rational evidence usually needs to be relatively short to speed up inference. To do so, reference systems are needed to determine whether something is long or short. For the rational evidence, the retrieved passages are commonly used as their reference system Wang et al. (2023); Zhao et al. (2024a); Zhu et al. (2024). On top of that, we use rational evidence as the reference system of the rationale.  They  can  be  formulated  as  follows:
    $R^{len}_{r}=\begin{cases}\frac{1}{1+e^{-(\mathrm{L}_{r}/\mathrm{L}_{e}-1)/\tau}},&\mathrm{L}_{r}\geq\mathrm{L}_{e},\\ \frac{1}{1+e^{-(1-\mathrm{L}_{e}/\mathrm{L}_{r})/\tau}},&\text{otherwise}.\end{cases}$ (6)
    $R^{len}_{e}=\begin{cases}1.0,&(1-\frac{\mathrm{L}_{e}}{\mathrm{L}_{P}})\geq\omega,\\ (1-\frac{\mathrm{L}_{e}}{\mathrm{L}_{P}})^{\gamma},&\text{otherwise}.\end{cases}$ (7)
    where $R^{len}_{r},R^{len}_{e}\in[0,1]$ denote the length rewards of rationale and rational evidence, respectively; $\mathrm{L}_{P}$, $\mathrm{L}_{r}$, and $\mathrm{L}_{e}$ denote the length of retrieved passages, the rationale, and the rational evidence, respectively; $\tau$ is a temperature coefficient; $\gamma\in[0,+\infty]$ is a hyperparameter to control the skewness of rewards; $\omega\in[0,1]$ is a threshold to avoid falling into a trivial solution. From Figure 4(a), we can see that the relatively long rationale ($x=2$) will be assigned large rewards, whereas a shorter rationale ($x=-2$) corresponds to small rewards. On the contrary, as shown in Figure 4(b), shorter rational evidence ($x=0.8$) will be assigned large rewards. By adjusting $\tau$ and $\gamma$, we can control the skewness or smoothness of length rewards, as shown in the line charts of different colors in Figure 4.
  • Format Reward. It focuses on whether the response obeys the defined format. Specifically, the model’s evidence reasoning, evidence extraction, and final output should be enclosed within the <reason></reason>, <extract></extract>, and <answer></answer> tags, respectively. Based on the above format requirement, the format reward  is  defined  as  follows:
    $R^{fmt}=\begin{cases}1.0,&\text{if correct format},\\ 0.0,&\text{if incorrect format}.\end{cases}$ (8)
Putting them together.   Having obtained the answer, length, and format rewards, we compute the final reward via a linear weighted sum, as complex weighting is not the focus of this work, and a linear one  generally  leads  to  satisfactory  performance:
$R^{final}=\alpha^{a}R^{ans}+\alpha^{l}R^{len}+\alpha^{f}R^{fmt},$ (9)
where $R^{ans}$ is the average answer reward of $R^{ans}_{r}$, $R^{ans}_{e}$ and $R^{ans}_{f}$; $R^{fmt}$ is the average length one of $R^{fmt}_{r}$ and $R^{fmt}_{e}$; $\alpha^{a}$, $\alpha^{l}$, and $\alpha^{f}$ are hyperparameters; $R^{final}$ is the final reward used in RL training.
3.3 Online Policy Optimization
Having obtained the final rewards of each response, we use the policy optimization algorithm GRPO [^4] Here, we adopt GRPO instead of Proximal Policy Optimization (PPO) Schulman et al. (2017), because PPO needs an  additional  critic  model  to  evaluate  policy  performance. Shao et al. (2024); DeepSeek-AI et al. (2025) to optimize LEAR, incentivizing rational evidence extraction capability in it via reinforcement learning. Specifically, for each input question $q$, GRPO first samples a group of responses $\mathcal{Y}=\{y_{1},y_{2},...,y_{G}\}$, where $G$ is the group size and each response consists of the rationale $r$, rational evidence $e$, and three outputs $o_{r}$, $o_{e}$, and $o_{f}$3.1). Subsequently, GRPO evaluates these responses based on the verifiable reward functions (§3.2) and obtains final rewards for each response =$\mathcal{R}=\{R_{1},R_{2},...,R_{G}\}$, where we omit the superscript “${final}$” for brevity. Different from PPO Schulman et al. (2017), GRPO directly compares the final rewards of candidate responses within the same group without needing an additional critic model. The advantage of the $i$-th response is determined through normalizing its reward $R_{i}$ using the mean and the standard deviation of rewards $\mathcal{R}$:
$A_{i}=\frac{R_{i}-\mathrm{mean}(\mathcal{R})}{\mathrm{std}(\mathcal{R})},$ (10)
where $\mathrm{mean}(\cdot)$ and $\mathrm{std}(\cdot)$ compute the average and standard deviation, respectively. However, GRPO’s group normalization may overly magnify the minor numerical fluctuation. Taking =$\mathcal{R}=\{0.49,0.51\}$ for example, the $\mathrm{mean}(R)$ is $0.5$, the $\mathrm{std}(R)$ is $0.01$, and the computed advantages are $\{-1.0,1.0\}$, which overly magnifies the minor numerical fluctuation. To mitigate this issue, we propose to clip $\mathrm{std}(R)$ to be at least $\epsilon_{std}$ to ensure that the denominator does not become too small:
$\tilde{A}_{i}=\frac{R_{i}-\mathrm{mean}(\mathcal{R})}{\mathrm{clip\_std}(\mathcal{R})},$ (11)
where $\mathrm{clip\_std}(\mathcal{R})=\mathrm{max}(\mathrm{std}(R),\epsilon_{std})$; $\epsilon_{std}$ is a hyperparameter. After obtaining refined advantages $\{\tilde{A}_{1},\tilde{A}_{2},...,\tilde{A}_{G}\}$, we can optimize the current policy model $\pi_{\theta}$ (i.e., the evidence extractor) by  maximizing the  following  objective  function:
$J_{\mathrm{GRPO}}(\theta)$ $=\mathbb{E}_{x\sim\mathcal{D},\,\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}(\cdot\mid x)}$ (12)
$\Biggl{[}\frac{1}{G}\sum_{i=1}^{G}\min\!\Bigl{(}\frac{\pi_{\theta}(\tilde{y}_{i}\mid x)}{\pi_{\theta_{\mathrm{old}}}(\tilde{y}_{i}\mid x)}\,\tilde{A}_{i},\,$
[…]
$-\,\beta\,D_{\mathrm{KL}}\bigl{(}\pi_{\theta}\,\big{\|}\,\pi_{\mathrm{ref}}\bigr{)}\Biggr{]},$
where $x=\{q,P\}$ denotes input samples drawn from the dataset $\mathcal{D}$; $y$ is the model’s response that consists of $\{r,e,o_{r},o_{e},o_{f}\}$, sampled from the old policy $\pi_{\theta_{\mathrm{old}}}$; $\tilde{y}$ is a self-contained response consisting of $\{r,e,o_{f}\}$, where $o_{f}$ is an output conditioned on full context, so they are self-contained and used for training; $\epsilon$ and $\beta$ are the PPO clipping hyperparameters and the weighting coefficient controlling the Kullback–Leibler (KL)-divergence, respectively; $\pi_{\mathrm{ref}}$ represents the reference policy.
4 Experiment
In this section, we conduct extensive experiments on three knowledge-intensive benchmark datasets to answer the following Research Questions (RQs): RQ1: How does rational evidence perform, w.r.t. downstream task performance, compared to other vanilla evidence? RQ2: How do the properties of rational evidence vary with the RL training process? RQ3: How does the inference efficiency of LEAR compare with that of different types of methods? RQ4: Can rational evidence perform robustly against retrieval noises? RQ5: How do different parts of the answer rewards contribute to the final model performance?
4.1 Experimental Setup
Datasets and Metrics.   We experiment on three knowledge-intensive benchmark QA datasets, i.e., Natural Questions (NQ) Kwiatkowski et al. (2019), TriviaQA (TQA) Joshi et al. (2017), and HotpotQA Yang et al. (2018), where the first two are open-domain question answering and the last one is multi-hop question answering. The detailed statistics of these three datasets are provided in Table 2, where the test set for HotpotQA is unavailable, and thus we use its dev set as a substitute for the test set. For evaluation, we adopt the Exact Match (EM) and unigram $\mathrm{F}_{1}$ score to evaluate QA performance. EM examines exact correctness while $\mathrm{F}_{1}$ calculates the degree of lexical overlap, offering a more fine-grained view of how well the prediction aligns with the golden answer. To measure the improvement of the computational efficiency, we employ the Compression Ratio (CR) following the previous works Hwang et al. (2024a); Pan et al. (2024), where CR computes the ratio of the total length of the retrieval passages to the length of the extracted evidence.
Dataset #Train #Dev #Test
NQ Kwiatkowski et al. (2019) 79.1k 8.7k 3.6k
TQA Joshi et al. (2017) 78.7k 8.8k 11.3k
HotoptQA Yang et al. (2018) 88.9k 5.6k 5.6k
Table 2: Statistics of the datasets.
Implementation Details.   We employ Qwen2.5-1.5B-Instruct and Qwen2.5-7B-Instruct [^5] https://github.com/QwenLM/Qwen2.5 Yang et al. (2025) as the initial models. We train these two models to extract rational evidence through full parameter fine-tuning for 1 epoch on the NQ dataset. The learning rate is $1e^{-6}$, and we select the best-performing model on the dev set. For modeling verifiable rewards, we tune $\tau$ and $\gamma$ within the ranges of $\{0.1,0.2,0.5,1.0\}$ and $\{0.1,0.3,0.5,0.8,1.0\}$, respectively. For the length threshold $\omega$ and the weighting coefficients (i.e., $\alpha^{a}$, $\alpha^{l}$, and $\alpha^{f}$), we empirically set $\omega$, $\alpha^{a}$, $\alpha^{l}$, and $\alpha^{f}$ as 0.9, 0.8, 0.1, 0.1, respectively, which commonly leads to satisfactory performance. For policy optimization, we set $\epsilon_{std}$ as 0.1 to stabilize training. During training, we set the PPO clipping hyperparameter $\epsilon$ and the coefficient $\beta$ controlling KL-divergence as 0.2 and $1e^{-2}$, respectively. For QA generators, we employ LoRA Yang et al. (2025) to fine-tune Qwen2.5-1.5B-Instruct and Qwen2.5-7B-Instruct to predict the golden answer $a$ based on the query and retrieved passages ($q$, $P$) or only the query (used for the baseline ‘Zero’), where LoRA rank is set as 8.
Dataset Recall NDCG
Train Dev Test Train Dev Test
NQ 78.74 73.07 74.07 68.30 61.95 63.08
TQA 82.35 77.97 77.77 76.06 70.32 70.35
HotoptQA 34.02 28.45 - 27.36 22.39 -
Table 3: Recall and NDCG of the top-5 retrieval passages in terms of training, development, and  test  sets.
Datasets Metrics WR VAR RAR
Zero Full SeleCtx LLMLingua-2 Recomp FilCo SEER CoT LEAR
1.5B size model
EM 13.74 41.97 24.68 29.36 37.40 36.62 36.93 37.70 41.14
F1 53.16 70.07 59.52 62.82 67.18 66.63 67.11 68.09 70.77
NQ CR - 1.0x 3.44x 4.51x 5.43x 16.3x 13.2x 4.56x 38.1x
EM 29.10 57.02 45.13 48.67 56.56 54.06 54.57 54.29 56.84
F1 64.62 80.13 73.40 75.29 79.85 78.56 78.81 79.56 80.85
TQA CR - 1.0x 3.38x 4.52x 5.35x 8.55x 10.3x 5.02x 38.8x
EM 12.36 19.20 15.30 16.64 18.52 18.18 18.60 19.52 20.46
F1 48.52 53.04 49.65 51.26 52.92 52.15 52.79 53.43 54.20
HotpotQA CR - 1.0x 3.40x 4.52x 5.44x 18.3x 15.5x 4.17x 33.0x
7B size model
EM 25.04 48.78 33.91 36.51 43.77 44.79 45.01 44.49 46.95
F1 60.88 74.40 65.71 67.77 71.61 72.30 72.76 73.09 73.99
NQ CR - 1.0x 3.44x 4.51x 5.43x 15.4x 12.4x 3.36x 17.9x
EM 47.31 65.34 58.72 60.50 64.40 63.46 64.20 63.91 64.76
F1 74.36 84.83 80.95 81.94 84.12 83.66 84.14 84.45 84.74
TQA CR - 1.0x 3.38x 4.52x 5.35x 7.77x 9.70x 3.33x 17.6x
EM 17.95 25.82 21.27 23.45 24.84 25.27 25.81 27.25 28.02
F1 53.07 58.50 54.90 56.65 58.02 58.15 58.63 59.84 60.30
HotpotQA CR - 1.0x 3.40x 4.52x 5.44x 17.5x 14.3x 3.32x 16.5x
[…]
Passage Retrieval.   Following previous works Wang et al. (2023); Zhu et al. (2024); Zhao et al. (2024a); Zhang et al. (2025a), we retrieve Wikipedia (Wiki) passages for all datasets. We use Dense Passage Retriever (DPR) Karpukhin et al. (2020) to retrieve the top-5 passages from all Wiki passages, where we use the December 2018 Wiki dump Karpukhin et al. (2020) as the retrieval corpus, and set the chunk size as 100 words. We use the retrieval toolkit, Tevatron Gao et al. (2023), to perform corpus encoding, query encoding, and passage retrieval. Table 3 shows the retrieval performance w.r.t. Recall@5 and NDCG@5, where the retrieval passage that contains the golden answer is treated as the positive passage. Besides, for HotpotQA, we compute Recall and NDCG only for the “bridge” questions, while ignoring yes/no and comparison questions, following previous works Jiang et al. (2024); Khalifa et al. (2023). The retrieval performance for HotpotQA is relatively low, as it is a multi-hop QA dataset, and answers are usually not spans in retrieval passages.
Comparison Baselines.   To verify the effectiveness of LEAR, we compare it with the following three groups of competitive baselines: (1) Without Refinement (WR) includes (i) Zero-shot (Zero) generates output relying on LLMs’ parametric knowledge; (ii) Full Passage (Full) feeds all retrieval passages into LLMs; (2) Vanilla Refinement (VAR) directly extracts evidence without explicit thinking, which includes (i) Select Context (SeleCtx) Li et al. (2023) identifies and prunes redundancy in the input context based on perplexity; (ii) LLMLingua-2 Pan et al. (2024) distills compression knowledge from GPT-4 to reduce crucial information losing; (iii) FilCo Wang et al. (2023) trains a context filtering model to identify key clues; (iv) BottleNeck Zhu et al. (2024) applies the information bottle theory to select SFT data used to optimize filtering; (v) SEER Zhao et al. (2024a) learns to extract desired evidence via self-aligned learning; (3) Rational Refinement (RAR) includes (i) Chain-of-Thought (CoT) Wei et al. (2022) generates query-related information from retrieval passages with explicit thinking.
(a) Answer reward dynamics on 1.5B model.
(b) Answer reward dynamics on 7B model.
(c) Length dynamics.
Figure 5: Training dynamics w.r.t. answer reward and response (including ‘reason’, ‘extract’, and ‘answer’) length.
4.2 Main Comparison (RQ1)
The overall comparison results on NQ, TQA, and HotpotQA are shown in Table 4. From the experimental results, we mainly have the following observations. (1) In all cases (18/18), LEAR achieves the best or second-best results, indicating the superiority of supplementing RAG with rational evidence. (2) It is surprising that directly employing LEAR trained on NQ and tested on OOD datasets (i.e., TQA and HotpotQA) yields such impressive performance. This demonstrates that online policy optimization endows LEAR with superior generalization. (3) Compared to ‘Full’, LEAR possesses an extremely high compression ratio (e.g., 38.1x with 1.5B size model on NQ) and its performance is very close to or even better than that of ‘Full’. Though the baseline models have relatively high compression ratios, they are accompanied by significant performance degradation. This demonstrates again that rational evidence benefits downstream performance than vanilla one and also improves the signal-to-noise ratio. (4) By comparing the 1.5B LEAR model with the 7B one, we find that the 7B one tends to extract more informative evidence than the 1.5B one. (5) LEAR considerably outperforms VAR methods in almost all cases and provides more compact evidence, demonstrating the necessity of explicit evidence reasoning. (6) In HotpotQA, RAR methods significantly outperform VAR ones, and even better than ‘Full’, indicating that rational refinement is important to multi-hop question answering. In conclusion, the above results and observations fully validate the effectiveness and efficiency of rational evidence.
4.3 Training Dynamics (RQ2)
The RL training dynamics regarding the answer rewards and response length are shown in Figure 5. Note that the generators used in Table 4 and Figure 5 are different. Therefore, the answer rewards in Figure 5 are slightly different from the $\mathrm{F_{1}}$ and EM results in Table 4. The results show that the answer rewards of full context ($o_{f}$), rationale ($o_{r}$), and rational evidence ($o_{e}$) are consistently improved during the process of reinforcement learning. And, it is not surprising that the answer reward of full context generally performs best. More specifically, on the 1.5B model, rationale and rational evidence perform very closely, while on the 7B model, rational evidence performs slightly worse than rationale. We think the main reason is that the 7B model can infer answers based on implicit cues, but rational evidence may compress some of them. As for response length dynamics, the response length of the 1.5B model decreases rapidly at the beginning of training and then increases slowly, while that of the 7B model decreases slowly all the time. We think the main reason is that a moderate response length is beneficial for improving answer rewards, because an overlong reasoning may confuse answer generation, and a too short reasoning may omit key cues. Therefore, as training continues, the response lengths of the 1.5B model and the 7B model tend to converge. In conclusion, the above observations provide some useful insights, e.g., elaborated answer and length rewards, for further research on the rational evidence extraction.
Models NQ TQA HotpotQA Avg
FilCo 0.64 0.82 0.59 0.68
CoT 0.55 0.70 0.71 0.65
LEAR 0.35 0.41 0.43 0.40
Table 5: Inference latency (seconds/query) on the 1.5B model, where the smaller the better.
4.4 Inference Efficiency (RQ3)
Table 5 presents the inference latency of LEAR compared with FilCo (a representative VAR method) and CoT (a representative RAR method) on the 1.5B model within a 1$\times$×A800 station, where the evaluation batch size and max new tokens are set as 64 and 768, respectively. It can be observed that the time required for evidence extraction using LEAR is considerably shorter than FilCo and CoT, facilitating effective application in online RAG systems. Additionally, it is surprising that the average inference latency of FilCo is slightly longer than that of CoT, while FilCo generates fewer new tokens on average. We find that the standard deviation in the lengths of the evidence generated by FilCo is significantly higher than that of COT. Specifically, the std of FilCo on NQ, TQA, and HotpotQA is 107.5, 136.1, and 81.7, while that of CoT is 67.9, 61.6, and 63.9. Due to the instability in the length of the generated evidence, FilCo takes longer in batching generation instead. In summary, the above results fully verify the efficiency of LEAR during deployment.
(a) NQ.
(b) TQA.
Figure 6: Performance comparisons w.r.t. data noise on the 1.5B model, where psgs is the abbr of ‘passages’.
Models Rationale Evidence Full Ctx Evidence+Rationale Evidence+Full Ctx Evidence+Rationale+Full Ctx
NQ 39.89 40.27 39.53 40.02 40.44 41.14
TQA 56.97 57.38 56.17 57.31 57.26 56.84
HotpotQA 20.30 19.73 20.13 20.34 20.23 20.46
Average 38.72 39.13 38.61 39.22 39.31 39.48
Table 6: Performance comparison (EM) of LEAR trained on different combinations of answer rewards.
4.5 Robustness Analysis (RQ4)
In real-world applications, RAG systems commonly suffer from data noise resulting from imperfect retrieval. To simulate this scenario, we randomly sample a certain number (i.e., 0, 2, 4, 6, and 8) of irrelevant passages for each test query, where each query is equipped with 5 retrieved relevant passages as well as sampled irrelevant passages. We experiment on the 1.5B model, and the experimental results are presented in Figure 6. The results show that adding noise considerably degrades the performance of FilCo, while the performance degradation of LEAR is relatively small, where the green line is always below the blue one. In particular, LEAR with 8 irrelevant psgs outperforms FilCo without noise. In conclusion, this fully verifies the robustness of LEAR against noise.
4.6 Ablation Study (RQ5)
To evaluate the impact of different answer rewards on model performance, we train 1.5B-sized LEAR models on different combinations of answer rewards with average weighting. As shown in Table 6, among models trained on a single answer reward, training with evidence answer reward $R^{ans}_{e}$ yields the highest average performance, demonstrating that the quality of evidence is the most important in optimizing rational evidence. For two-answer reward combinations, ‘Evidence+Full Ctx’ slightly outperforms ‘Evidence+Rationale’, with noticeable gains on NQ in particular, indicating that the optimization of ‘Full Ctx’ matters. Notably, incorporating all three answer rewards yields the best performance on NQ and HotpotQA, as well as the highest average score, highlighting the benefits of multi-faceted reward assessment. In conclusion, the above results clearly manifest the contribution and necessity of each answer reward.
5 Related Works
5.1 Retrieval-Augmented Generation
Retrieval-Augmented Generation (RAG) has been prevalent in LLMs, which effectively enhances LLMs with external non-parametric knowledge to remedy their outdated, incorrect, or incomplete internal parametric knowledge Lewis et al. (2020); Wu et al. (2022); Guu et al. (2020); Izacard et al. (2023); Asai et al. (2024). The pioneering attempts Lewis et al. (2020); Guu et al. (2020) demonstrated that augmenting the input context of LLMs with retrieved passages yields significant improvements, especially in knowledge-intensive tasks. While prior work usually retrieves straightforwardly at the beginning, dynamic RAG Asai et al. (2024); Jeong et al. (2024); Jiang et al. (2023) has been proposed to adaptively retrieve passages based on the demand of generation and the complexity of the query. Recently, agentic RAG Zhang et al. (2025b) has been proposed to interleave retrieval and reasoning to tackle complex queries. In particular, this kind of work Song et al. (2025); Jin et al. (2025); Trivedi et al. (2023) focuses on prompting or fine-tuning LLMs as search agents that interact with search tools to get external knowledge autonomously and on demand through direct prompting/reinforcement learning.
5.2 RAG with Context Compression
RAG systems usually concatenate all retrieved passages as the context of LLMs. However, this may introduce data noise and computational overhead due to imperfect retrieval and overlong context. Recently, many works have attempted to compress context and retain important information, mainly including two categories: (1) Reranking methods, which rerank retrieved passages and retain top-ranked passages Hwang et al. (2024b); Mortaheb et al. (2025); Mao et al. (2021) (2) Summarization methods, which extract relevant information from retrieved passages and consolidate them into a narrativeWang et al. (2023); Zhao et al. (2024a); Zhu et al. (2024); Xu et al. (2024). Despite effectiveness, existing works neglect explicit evidence reasoning, which risks filtering out key clues as well as struggles with generalization.
6 Conclusion
In this paper, we first unveil the limitations of the vanilla evidence extraction paradigm and explore the potential of the rational evidence extraction paradigm to solve these limitations. Specifically, we propose LEAR, a rational evidence extraction reinforcement learning framework. In particular, we unify evidence reasoning and extraction into one unified response, and devise verifiable reward functions to guide the optimization of LEAR for compact and high-quality evidence, where knowledge token masking and std clipping are applied to avoid information leakage and stabilize RL training, respectively. Extensive experiments fully demonstrate the superiority of LEAR w.r.t. performance, generalization, and robustness.
Limitations
Despite our innovations and improvements, it does have limitations, especially cascaded generation between the rationale and the rational evidence. This indicates that the rationale must be generated before the rational evidence, which increases the inference latency to some extent. Although the results in Table 5 show that the inference efficiency of LEAR remains superior to that of the baseline, we believe that further improvements in reasoning efficiency are necessary to achieve even higher inference speed. For example, reasoning should be conducted only when necessary Jiang et al. (2025), i.e., generating rational evidence directly for easy instances. We leave it for future research.
Appendix A Prompts
We provide the prompts that are used for retrieval-augmented QA, closed-book QA, CoT generation, and rational evidence extraction in Table 7, Table 8, Table 9, and Table 10, respectively.
Appendix B Case Study
Through extensive case studies, we found the reasons for the failure of vanilla evidence mainly lie in threefold, i.e., Incompleteness, Irrelevance, and Inaccuracy, termed as “3I”. For Incompleteness, an example on the left of Figure 7, vanilla evidence provides information about the award category (Best Actor) but fails to provide the specific name of the award winner. For Irrelevance, an example in the middle of Figure 7, vanilla evidence provides irrelevant information about a different work and year. For Inaccuracy, an example in the right of Figure 7, vanilla evidence contains a factual error, attributing the opera "The Midsummer Marriage" to the wrong composer (Benjamin Britten instead of Michael Tippett). On the contrary, rational evidence demonstrates a greater capacity for understanding the query’s intent, verifying factual accuracy, and reasoning to arrive at the correct and relevant evidence. It moves beyond simple evidence extraction towards reasonable evidence extraction.
Prompt for Retrieval-Augmented QA [Instruction]
You are a helpful assistant. Your task is:
1. Read the given question and use the documents provided to answer the question. 2. If the documents don’t work, please answer the question based on your own knowledge. Question: {question} Document: {document} Answer:
Table 7: The prompt for retrieval-augmented QA.
Prompt for Closed-book QA [Instruction]
You are a helpful assistant. Your task is:
1. Read the given question and then answer the question directly. 2. Give a short answer to the question based on your own knowledge. Question: {question} Answer:
Table 8: The prompt for closed-book QA.
Prompt for CoT Generation [Instruction]
You are a helpful assistant. Your task is:
Read the given documents, and answer the question below. Question: {question} Document: {document} Let’s think step by step.
Table 9: The prompt for Chain-of-Thought generation.
Prompt for Rational Evidence Extraction [Instruction]
You are a highly skilled knowledge reasoner and extractor.
Your task is to carefully read the given question and passages to reason how the passages lead to the answer and extract relevant information that may be used to answer the question. Follow these steps: 1. In the <reason></reason> tag, perform the following steps. Question Analysis: Analyze the question to understand the specific information they are seeking. Identify the key concepts, entities, and relationships involved. Passage Analysis: For each passage, carefully read and identify sentences or phrases that are useful for answering the given question. 2. In the <extract></extract> tag, synthesize useful information from the passages into a coherent narrative. Organize the information logically and concisely. 3. In <answer></answer> tags, give a short answer to the given question, based on the passages, reasoning information, and extracted knowledge. If none of them work, please answer the question based on your knowledge. Question: {question} Passages: {passages}
Table 10: The prompt for rational evidence extraction, where generation is terminated when encountering token ‘</extract>’.
Figure 7: Three main failure issues, termed as 3I: (1) Incompleteness, where the evidence provides some relevant information, but lacks the key clues needed to answer the questions; (2) Irrelevance, where the evidence provides information about the wrong entity, event, or topic; and (3) Inaccuracy, where the evidence contains incorrect information, such as wrong dates, names, or relationships.

### v3

@@
 (2025); Mao et al.
+(2021); Chen et al.
 (2025); and (2) Summarizing retrieval contents into a coherent narrative with higher signal to noise ratio Wang et al.
@@
 (2024a).
-They typically create the training data using hand-crafted strategies, e.g., String Inclusion that measures whether the golden answer is included in the candidate passage, and Lexical Overlap, calculating the unigram overlap between the golden answer and candidate passage.
+They typically create the training data using hand-crafted strategies, e.g., String Inclusion that measures whether the golden answer is included in the candidate passage, and Lexical Overlap that calculates the unigram overlap between the golden answer and candidate passage.
 Despite effectiveness, existing methods re-rank or summarize retrieval contents straightforwardly without explicit thinking, which risks filtering out key clues due to a lack of a deep understanding of retrieved contents.
@@
 2.1 Problem Statement
-In RAG Lewis et al.
-(2020); Wu et al.
-(2022), LLMs are given an input query $q$ and a set of top-$k$ retrieved passages $P=\{p_{1},p_{2},...,p_{k}\}$, which aim to generate an output $o$ closing to the golden answer $a$, conditioned on ($q$, $P$).
-In the traditional RAG paradigm, top-$k$ retrieved passages are directly fed into LLMs.
-However, retrieved passages $P$ often contain irrelevant or noisy contents Wang et al.
-(2023); Zhao et al.
-(2024a), which considerably degrade LLMs’ generation quality.
-Given that, an additional evidence extraction model is proposed to condense $P$ into a concise, query-relevant context $e$ to improve generation quality and speed.
-Formally, the paradigm of RAG with evidence extraction can be formulated as follows:
+[…]
 $e=\mathcal{M}_{\mathcal{E}}(\cdot|q,P),\quad o=\mathcal{M}_{\mathcal{G}}(\cdot|q,e),$ (1)
@@
 $s.t.\ |e|\ll|P|$ (3)
-where $\mathcal{U}(\cdot)$ is a utility function (e.g., $\mathrm{F}_{1}$ score) and is used to measure the quality of outputs conditioned on the golden answer $a$; $|e|$ and $|P|$ denote the length of evidence  and  passages, respectively.
+[…]
 (a) Results on NQ.
@@
 $R^{final}=\alpha^{a}R^{ans}+\alpha^{l}R^{len}+\alpha^{f}R^{fmt},$ (9)
-where $R^{ans}$ is the average answer reward of $R^{ans}_{r}$, $R^{ans}_{e}$ and $R^{ans}_{f}$; $R^{fmt}$ is the average length one of $R^{fmt}_{r}$ and $R^{fmt}_{e}$; $\alpha^{a}$, $\alpha^{l}$, and $\alpha^{f}$ are hyperparameters; $R^{final}$ is the final reward used in RL training.
+where $R^{ans}$ is the average answer reward of $R^{ans}_{r}$, $R^{ans}_{e}$ and $R^{ans}_{f}$; $R^{len}$ is the average length one of $R^{len}_{r}$ and $R^{len}_{e}$; $\alpha^{a}$, $\alpha^{l}$, and $\alpha^{f}$ are hyperparameters; $R^{final}$ is the final reward used in RL training.
 3.3 Online Policy Optimization
@@
 However, GRPO’s group normalization may overly magnify the minor numerical fluctuation.
-Taking =$\mathcal{R}=\{0.49,0.51\}$ for example, the $\mathrm{mean}(R)$ is $0.5$, the $\mathrm{std}(R)$ is $0.01$, and the computed advantages are $\{-1.0,1.0\}$, which overly magnifies the minor numerical fluctuation.
+Taking =$\mathcal{R}=\{0.49,0.51\}$ for example, the $\mathrm{mean}(R)$ is 0.50.50.5, the $\mathrm{std}(R)$ is 0.010.010.01, and the computed advantages are {$,1.0\}$ 1.0 }, which overly magnifies the minor numerical fluctuation.
 To mitigate this issue, we propose to clip $\mathrm{std}(R)$ to be at least $\epsilon_{std}$ to ensure that the denominator does not become too small:
@@
 Table 2: Statistics of the datasets.
-Implementation Details.
-We employ Qwen2.5-1.5B-Instruct and Qwen2.5-7B-Instruct [^5] https://github.com/QwenLM/Qwen2.5 Yang et al.
-(2025) as the initial models.
-We train these two models to extract rational evidence through full parameter fine-tuning for 1 epoch on the NQ dataset.
-The learning rate is $1e^{-6}$, and we select the best-performing model on the dev set.
-For modeling verifiable rewards, we tune $\tau$ and $\gamma$ within the ranges of $\{0.1,0.2,0.5,1.0\}$ and $\{0.1,0.3,0.5,0.8,1.0\}$, respectively.
-For the length threshold $\omega$ and the weighting coefficients (i.e., $\alpha^{a}$, $\alpha^{l}$, and $\alpha^{f}$), we empirically set $\omega$, $\alpha^{a}$, $\alpha^{l}$, and $\alpha^{f}$ as 0.9, 0.8, 0.1, 0.1, respectively, which commonly leads to satisfactory performance.
-For policy optimization, we set $\epsilon_{std}$ as 0.1 to stabilize training.
-During training, we set the PPO clipping hyperparameter $\epsilon$ and the coefficient $\beta$ controlling KL-divergence as 0.2 and $1e^{-2}$, respectively.
-For QA generators, we employ LoRA Yang et al.
-(2025) to fine-tune Qwen2.5-1.5B-Instruct and Qwen2.5-7B-Instruct to predict the golden answer $a$ based on the query and retrieved passages ($q$, $P$) or only the query (used for the baseline ‘Zero’), where LoRA rank is set as 8.
+[…]
 Dataset Recall NDCG
@@
 (2025); Mao et al.
+(2021); Chen et al.
 (2025) (2) Summarization methods, which extract relevant information from retrieved passages and consolidate them into a narrative Wang et al.
@@
 (2024).
-Despite effectiveness, existing works neglect explicit evidence reasoning, which risks filtering out key clues as well as struggles with generalization.
+While effective, existing works neglect evidence reasoning, which risks filtering out key clues and struggles with generalization.
 6 Conclusion

### v4

@@
 Previous methods extract evidence straightforwardly without explicit thinking, which risks filtering out key clues and struggles with generalization.
-To this end, we propose LEAR, which learns to extract rational evidence by (1) explicitly reasoning to identify potential cues within retrieval contents first, and then (2) consciously extracting to avoid omitting any key cues helpful for answering questions.
-Specifically, we frame evidence reasoning and extraction into one unified response for end-to-end training; apply knowledge token masks for disentanglement to derive reasoning-based and extraction-based answers; and devise three types of verifiable reward functions, including answer, length, and format, to update the model via the policy optimization algorithm.
-Extensive experiments on three benchmark datasets show the effectiveness of LEAR, providing compact and high-quality evidence, improving the accuracy of downstream tasks, and promoting effective application in online RAG systems [^1] The code, data, and models will be available at https://github.com/HITsz-TMG/LEAR..
+To this end, we propose EviOmni, which learns to extract rational evidence by (1) explicitly reasoning to identify potential cues within retrieval contents first, and then (2) consciously extracting to avoid omitting any key cues helpful for answering questions.
+Specifically, we frame evidence reasoning and evidence extraction into one unified response for end-to-end training; apply knowledge token masks for disentanglement to derive reasoning-based and extraction-based answers; and devise three types of verifiable reward functions, including answer, length, and format, to update the model via the policy optimization algorithm.
+Extensive experiments on three benchmark datasets show the effectiveness of EviOmni, providing compact and high-quality evidence, improving the accuracy of downstream tasks, and promoting effective application in online RAG systems [^1] The code, data, and models will be available at https://github.com/HITsz-TMG/EviOmni..
 Figure 1: Motivating example, where key clues are marked in green: 1 The key clues are filtered out, and LLMs answer incorrectly; 2 The key clues are extracted successfully, guided by the evidence reasoning.
@@
 This motivates us to develop a rational evidence extractor for RAG, which deeply reasons and then consciously extracts as shown in the bottom layer of Figure 1.
-In this work, we propose LEAR, which LEArns to extract Rational evidence via reinforcement learning for RAG.
-Specifically, LEAR frames evidence reasoning (enclosed within <reason> and </reason>) and extraction (enclosed within <extract> and </extract>) into one response for end-to-end training and generates multiple responses to collect trajectories that contain good and bad experiences from the policy.
-After that, LEAR applies two knowledge token masks on each response to disentangle evidence reasoning and extraction for deriving reasoning-based and extraction-based answers (enclosed within <answer> and </answer> tags) to assess their quality respectively.
+In this work, we propose EviOmni, which learns to extract rational evidence via reinforcement learning for RAG.
+Specifically, EviOmni frames evidence reasoning (enclosed within <reason> and </reason>) and extraction (enclosed within <extract> and </extract>) into one response for end-to-end training and generates multiple responses to collect trajectories that contain good and bad experiences from the policy.
+After that, EviOmni applies two knowledge token masks on each response to disentangle evidence reasoning and extraction for deriving reasoning-based and extraction-based answers (enclosed within <answer> and </answer> tags) to assess their quality respectively.
 Finally, we devise three types of rule-based verifiable reward functions, including answer, length, and format, to guide model optimization via Group Relative Policy Optimization (GRPO) Shao et al.
 (2024).
-As such, LEAR extracts rational evidence by reasoning about retrieval contents and then consolidating key cues into a concise and useful narrative, which is fed into LLMs to generate more accurate output.
+As such, EviOmni extracts rational evidence by reasoning about retrieval contents and then consolidating key cues into a concise and useful narrative, which is fed into LLMs to generate more accurate output.
 Our main contributions can be summarized as follows:
   • We unveil the major issue that hinders evidence extraction for RAG, i.e., lacking a deep contextual understanding of retrieval contents, which risks filtering out key clues.
-  • We propose a novel rational evidence extraction learning framework, LEAR, which learns to discern potential cues in retrieval contents and then consciously consolidates any key cues into a concise yet very helpful evidence.
-  • We conduct extensive experiments on three benchmark datasets, where the results fully demonstrate the superiority of LEAR in terms of performance, generalization, efficiency, as well as robustness against retrieval noise.
+  • We propose a novel rational evidence extraction learning framework, EviOmni, learning to discern potential cues in retrieval contents and consciously consolidates any key cues into a concise yet very helpful evidence.
+  • We conduct extensive experiments on three benchmark datasets, where the results fully demonstrate the superiority of EviOmni in terms of performance, generalization, efficiency, and robustness against retrieval noise.
 2 Preliminaries
@@
 7.30% on HotpotQA, fully demonstrating the superiority of the rational evidence.
-Figure 3: The overall system framework of the proposed LEAR and previous vanilla paradigm.
+Figure 3: The overall system framework of the proposed EviOmni and previous vanilla paradigm.
 (a) Compared to vanilla evidence, our rational evidence is more skilled in extracting key cues.
@@
 (2023).
-(c) The proposed framework LEAR incentivizes rational evidence extraction capability in the extractor $\pi_{\theta}$ via reinforcement learning.
+(c) The proposed framework EviOmni incentivizes rational evidence extraction capability in the extractor $\pi_{\theta}$ through RL.
 We also provide a detailed failure analysis of vanilla evidence in Appendix B.
@@
 3 Methodology
-The overall framework of LEAR is illustrated in Figure 3.
+The overall framework of EviOmni is illustrated in Figure 3.
 In this section, we first introduce the acquisition of rational evidence (§3.1).
@@
 (2025), which guides the optimization process towards desired properties.
-To train LEAR via RL, we design three types of rule-based verifiable reward functions w.r.t.
+To train EviOmni via RL, we design three types of rule-based verifiable reward functions w.r.t.
 three primary desired properties for evidence extraction: (1) The correctness of the generated outputs; (2) The comprehensiveness of rationale and the conciseness of rational evidence; and (3) The format specification of the responses.
@@
 (2024); DeepSeek-AI et al.
-(2025) to optimize LEAR, incentivizing rational evidence extraction capability in it via reinforcement learning.
+(2025) to optimize EviOmni, incentivizing rational evidence extraction capability in it via reinforcement learning.
 Specifically, for each input question $q$, GRPO first samples a group of responses $\mathcal{Y}=\{y_{1},y_{2},...,y_{G}\}$, where $G$ is the group size and each response consists of the rationale $r$, rational evidence $e$, and three outputs $o_{r}$, $o_{e}$, and $o_{f}$3.1).
@@
 RQ2: How do the properties of rational evidence vary with the RL training process?
-RQ3: How does the inference efficiency of LEAR compare with that of different types of methods?
+RQ3: How does the inference efficiency of EviOmni compare with that of different types of methods?
 RQ4: Can rational evidence perform robustly against retrieval noises?
@@
 Datasets Metrics WR VAR RAR
-Zero Full SeleCtx LLMLingua-2 Recomp FilCo SEER CoT LEAR
+Zero Full SeleCtx LLMLingua-2 Recomp FilCo SEER CoT EviOmni
 1.5B size model
@@
 Comparison Baselines.
-To verify the effectiveness of LEAR, we compare it with the following three groups of competitive baselines: (1) Without Refinement (WR) includes (i) Zero-shot (Zero) generates output relying on LLMs’ parametric knowledge; (ii) Full Passage (Full) feeds all retrieval passages into LLMs; (2) Vanilla Refinement (VAR) directly extracts evidence without explicit thinking, which includes (i) Select Context (SeleCtx) Li et al.
+To verify the effectiveness of EviOmni, we compare it with the following three groups of competitive baselines: (1) Without Refinement (WR) includes (i) Zero-shot (Zero) generates output relying on LLMs’ parametric knowledge; (ii) Full Passage (Full) feeds all retrieval passages into LLMs; (2) Vanilla Refinement (VAR) directly extracts evidence without explicit thinking, which includes (i) Select Context (SeleCtx) Li et al.
 (2023) identifies and prunes redundancy in the input context based on perplexity; (ii) LLMLingua-2 Pan et al.
@@
 From the experimental results, we mainly have the following observations.
-(1) In all cases (18/18), LEAR achieves the best or second-best results, indicating the superiority of supplementing RAG with rational evidence.
-(2) It is surprising that directly employing LEAR trained on NQ and tested on OOD datasets (i.e., TQA and HotpotQA) yields such impressive performance.
-This demonstrates that online policy optimization endows LEAR with superior generalization.
-(3) Compared to ‘Full’, LEAR possesses an extremely high compression ratio (e.g., 38.1x with 1.5B size model on NQ) and its performance is very close to or even better than that of ‘Full’.
+(1) In all cases (18/18), EviOmni achieves the best or second-best results, indicating the superiority of supplementing RAG with rational evidence.
+(2) It is surprising that directly employing EviOmni trained on NQ and tested on OOD datasets (i.e., TQA and HotpotQA) yields such impressive performance.
+This demonstrates that online policy optimization endows EviOmni with superior generalization.
+(3) Compared to ‘Full’, EviOmni possesses an extremely high compression ratio (e.g., 38.1x with 1.5B size model on NQ) and its performance is very close to or even better than that of ‘Full’.
 Though the baseline models have relatively high compression ratios, they are accompanied by significant performance degradation.
 This demonstrates again that rational evidence benefits downstream performance than vanilla one and also improves the signal-to-noise ratio.
-(4) By comparing the 1.5B LEAR model with the 7B one, we find that the 7B one tends to extract more informative evidence than the 1.5B one.
-(5) LEAR considerably outperforms VAR methods in almost all cases and provides more compact evidence, demonstrating the necessity of explicit evidence reasoning.
+(4) By comparing the 1.5B EviOmni model with the 7B one, we find that the 7B one tends to extract more informative evidence than the 1.5B one.
+(5) EviOmni considerably outperforms VAR methods in almost all cases and provides more compact evidence, demonstrating the necessity of explicit evidence reasoning.
 (6) In HotpotQA, RAR methods significantly outperform VAR ones, and even better than ‘Full’, indicating that rational refinement is important to multi-hop question answering.
@@
 CoT 0.55 0.70 0.71 0.65
-LEAR 0.35 0.41 0.43 0.40
+EviOmni 0.35 0.41 0.43 0.40
 Table 5: Inference latency (seconds/query) on the 1.5B model, where the smaller the better.
 4.4 Inference Efficiency (RQ3)
-Table 5 presents the inference latency of LEAR compared with FilCo (a representative VAR method) and CoT (a representative RAR method) on the 1.5B model within a 1$\times$×A800 station, where the evaluation batch size and max new tokens are set as 64 and 768, respectively.
-It can be observed that the time required for evidence extraction using LEAR is considerably shorter than FilCo and CoT, facilitating effective application in online RAG systems.
-Additionally, it is surprising that the average inference latency of FilCo is slightly longer than that of CoT, while FilCo generates fewer new tokens on average.
+Table 5 presents the inference latency of EviOmni compared with FilCo (a representative VAR method) and CoT (a representative RAR method) on the 1.5B model within a 1$\times$×A800 station, where the evaluation batch size and max new tokens are set as 64 and 768, respectively.
+It can be observed that the time required for evidence extraction using EviOmni is considerably shorter than FilCo and CoT, facilitating effective application in online RAG systems.
+It is surprising that the average inference latency of FilCo is slightly longer than that of CoT, while FilCo generates fewer new tokens on average.
 We find that the standard deviation in the lengths of the evidence generated by FilCo is significantly higher than that of COT.
@@
 Due to the instability in the length of the generated evidence, FilCo takes longer in batching generation instead.
-In summary, the above results fully verify the efficiency of LEAR during deployment.
+In all, the above results fully verify the efficiency of EviOmni during deployment.
 (a) NQ.
@@
 HotpotQA 20.30 19.73 20.13 20.34 20.23 20.46
-Average 38.72 39.13 38.61 39.22 39.31 39.48
-Table 6: Performance comparison (EM) of LEAR trained on different combinations of answer rewards.
+Average 39.05 39.13 38.61 39.22 39.31 39.48
+Table 6: Performance comparison (EM) of EviOmni trained on different combinations of answer rewards.
 4.5 Robustness Analysis (RQ4)
 In real-world applications, RAG systems commonly suffer from data noise resulting from imperfect retrieval.
-To simulate this scenario, we randomly sample a certain number (i.e., 0, 2, 4, 6, and 8) of irrelevant passages for each test query, where each query is equipped with 5 retrieved relevant passages as well as sampled irrelevant passages.
+To simulate this scenario, we randomly sample a certain number (i.e., 0, 2, 4, 6, and 8) of irrelevant passages for each test query, where each query is equipped with 5 retrieved relevant passages and sampled irrelevant passages.
 We experiment on the 1.5B model, and the experimental results are presented in Figure 6.
-The results show that adding noise considerably degrades the performance of FilCo, while the performance degradation of LEAR is relatively small, where the green line is always below the blue one.
-In particular, LEAR with 8 irrelevant psgs outperforms FilCo without noise.
-In conclusion, this fully verifies the robustness of LEAR against noise.
+The results show that adding noise considerably degrades the performance of FilCo, while the performance degradation of EviOmni is relatively small, where the green line is always below the blue one.
+In particular, EviOmni with 8 irrelevant psgs outperforms FilCo without noise.
+In all, this fully verifies the robustness of EviOmni to noise.
 4.6 Ablation Study (RQ5)
-To evaluate the impact of different answer rewards on model performance, we train 1.5B-sized LEAR models on different combinations of answer rewards with average weighting.
+To evaluate the impact of different answer rewards on model performance, we train 1.5B-sized EviOmni models on different combinations of answer rewards with average weighting.
 As shown in Table 6, among models trained on a single answer reward, training with evidence answer reward $R^{ans}_{e}$ yields the highest average performance, demonstrating that the quality of evidence is the most important in optimizing rational evidence.
-For two-answer reward combinations, ‘Evidence+Full Ctx’ slightly outperforms ‘Evidence+Rationale’, with noticeable gains on NQ in particular, indicating that the optimization of ‘Full Ctx’ matters.
-Notably, incorporating all three answer rewards yields the best performance on NQ and HotpotQA, as well as the highest average score, highlighting the benefits of multi-faceted reward assessment.
-In conclusion, the above results clearly manifest the contribution and necessity of each answer reward.
+For two-answer reward combinations, ‘Evidence+Full Ctx’ slightly outperforms ‘Evidence+Rationale’, with significant gains on NQ, indicating that the optimization of ‘Full Ctx’ matters.
+Notably, incorporating all three answer rewards yields the best performance on NQ and HotpotQA, and the highest average score, indicating the benefits of multi-faceted reward assessment.
+In all, the above results manifest the contribution and necessity of each answer reward.
 5 Related Works
@@
 In this paper, we first unveil the limitations of the vanilla evidence extraction paradigm and explore the potential of the rational evidence extraction paradigm to solve these limitations.
-Specifically, we propose LEAR, a rational evidence extraction reinforcement learning framework.
-In particular, we unify evidence reasoning and extraction into one unified response, and devise verifiable reward functions to guide the optimization of LEAR for compact and high-quality evidence, where knowledge token masking and std clipping are applied to avoid information leakage and stabilize RL training, respectively.
-Extensive experiments fully demonstrate the superiority of LEAR w.r.t.
+Specifically, we propose EviOmni, a rational evidence extraction reinforcement learning framework.
+In particular, we unify evidence reasoning and extraction into one unified response, and devise verifiable reward functions to guide the optimization of EviOmni for compact and high-quality evidence, where knowledge token masking and std clipping are applied to avoid information leakage and stabilize RL training, respectively.
+Extensive experiments fully manifest the superiority of EviOmni w.r.t.
 performance, generalization, and robustness.
@@
 This indicates that the rationale must be generated before the rational evidence, which increases the inference latency to some extent.
-Although the results in Table 5 show that the inference efficiency of LEAR remains superior to that of the baseline, we believe that further improvements in reasoning efficiency are necessary to achieve even higher inference speed.
+Although the results in Table 5 show that the inference efficiency of EviOmni remains superior to that of the baseline, we believe that further improvements in reasoning efficiency are necessary to achieve even higher inference speed.
 For example, reasoning should be conducted only when necessary Jiang et al.